The point (-2,2) is a solution to the following linear system. Find the values for p and q.
-5px+(7q-1)y=-4
(p+10)x-(4q+1)y=-32

Answer by ewatrrr(24785)


Hi,
1)Substituting (-2,2)in the first equation
-5p(-2)+(7q-1)*2=-4
10p+ 14q-2 =-4
10p+ 14q = -2
2)Substituting (-2,2)in the second equation
(p+10)(-2)-(4q+1)*2=-32
-20p - 8q - 2 = 32
-20p - 8q = 34
Multiplying first result by 2 and adding together
20p+ 28q = -4
-20p - 8q = 34
20q = 30
q = 3/2
10p + 21 = -2
p = -23/10
checking our answer
10*(-23/10) +14(3/2) = -23 + 21 = -2